OpenStudy (anonymous):

find the length of a rectangular lot with a perimeter of a 82 m if the length is 5 m more than the width. a=46m b=23m c=41m d=18m

OpenStudy (anonymous):

l = 5 + w w + w +5 = 82/2 => 2w + 5 = 41 => 2w = 36 => w = 18 l = 18+5